OpenStudy (anonymous):

What are the zeros of the function? What are the multiplicities? \[(fx)=5x^3-5x^2-30x\] Factor: \[(5x)(x^2 - x - 6)\] What do I do from there?

hartnn (hartnn):

to get the zeros, equate f(x) = 0 5x (x^2-x-6) = 0 5x = 0 or x^2 -x-6 =0 can you solve that quadratic ?

OpenStudy (anonymous):

\[x\; =\; \frac {-b\; \pm\; \sqrt {b^2\; -\; 4ac}}{2a}\\\] How would I in put that into the formula?

hartnn (hartnn):

you only know quadratic formula method ? do u also know factor method to solve quadratics ?

OpenStudy (anonymous):

Nope :/

hartnn (hartnn):

ok, lets do it by quadratic formula :)